# On the Non-Identifiability of Steering Vectors in Large Language Models

### v2

@@ 6.4 Robustness of Non-Identifiability Under Distribution Shift
We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability. We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing), and safety-style (refusal patterns, ethical dilemmas). For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
Table 2 and Figure 4 present Cohen’s $d$ effect sizes across all trait-environment combinations for both models. Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$). Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across extraction contexts. ID = in-distribution; Top = topic shift; Gen = genre shift; Saf = safety-style prompts.
Model Trait ID Top Gen Saf
@@
Figure 4: Multi-environment validation results. Environment-dependent effect sizes and cross-environment matrices show small-to-moderate Cohen’s $d$ values across traits and prompt contexts, confirming robustness of observational equivalence under distributional shift.
Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models). Politeness exhibits higher variability ($d$ reaching 0.52–0.63 in genre-shift and safety contexts) but remains within the small-effect range. Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama). Figure 4 visualizes these patterns through environment-dependent effect sizes (left panels) and cross-environment transfer matrices (right panels), showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
Implications. The persistence of small effect sizes across diverse prompt distributions demonstrates that non-identifiability is not an artifact of in-distribution prompts but a fundamental geometric property. Critically, this shows that naive distributional diversity without explicit invariance constraints or structural assumptions is insufficient to recover identifiable representations.
@@ 7 Conclusion
We have established that steering vectors are fundamentally non-identifiable under standard observational regimes, with infinitely many distinct directions producing equivalent behavioral effects. This non-identifiability persists across models, traits, and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may not yield interpretable representations. Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone.
8 Limitations
Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor). While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions. Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human judgments, providing a conservative test of observational equivalence.
Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions, but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization, or invariance objectives (IRM) can recover identifiable representations. While theoretical frameworks suggest these approaches might enable identifiability, empirical validation across multiple models and experimental designs remains an important direction for future work. Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
9 Future Work
Our results establish the necessity of validation beyond input-output behavior. Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components, and characterizing how null-space dimensionality scales with model size.
Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability. Human evaluation would provide stronger validation than our lexical heuristics, and direct measurement of null-space dimensionality would quantify our theoretical predictions.
10 Acknowledgments
We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments. We also acknowledge the HuggingFace platform and maintainers of the open-source models (Qwen2.5-3B-Instruct, Llama-3.1-8B-Instruct) used in this work.
@@ Appendix A Logit-Level Equivalence Test
Values $R<1$ indicate that orthogonal perturbations induce smaller logit deviations than random directions. In all tested models and traits, this condition holds. Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions. The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$, and top-10 token overlap ranges from $86\%$ to $96\%$.
These results provide direct evidence that equivalence is not an artifact of semantic measurement. While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes. Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement, and top-$k$ overlap for all models and traits.
Table 3: Logit-level equivalence statistics.
Model Trait $R$ (v_perp / random) Token Agreement Top-10 Overlap
@@
Qwen-3B Formality 0.76 94% 96%
Qwen-3B Politeness 0.75 93% 96%
@@
Llama-8B Formality 0.49 93% 95%
Llama-8B Politeness 0.65 91% 94%
Appendix B Prompt Construction Details
B.1 Contrastive Prompt Pairs
Trait Positive ($x^{+}$) Negative ($x^{-}$)
Table 4: Example contrastive prompt pairs for each semantic trait.
B.2 Steering Extraction and Evaluation Pipeline
Figure 5: Steering extraction and evaluation pipeline.
Appendix C Detailed Proof of Proposition 1
We provide proofs for establishing non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
C.1 Null-Space Ambiguity (Constructive Proof)
Setup. Consider the local linear approximation of the steering effect:
@@ C.1 Null-Space Ambiguity (Constructive Proof)
Step 1: Null space characterization.
@@
Step 2: Rank Bound.
@@
In modern language models:
  • Hidden dimension: $d\approx 4000$ (typical)
  • Vocabulary size: $V\approx 50000$ (typical)
  • Therefore: $\max\text{rank}(J_{\ell})=d$
Step 3: Effective rank is much lower.
@@
Intuition: In overparameterized LLMs, the effective rank is expected to be strictly less than $d$ due to the low-dimensional structure of output distributions. Therefore $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is generically positive but this is not required for the proof—the argument establishes non-identifiability whenever $\dim(\mathcal{N})\geq 1$.
Step 4: Constructing equivalent vectors.
@@
$o(x,v^{\prime},\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}=o(x,0,0)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha).$ (18)
Step 5: Infinitely many distinct solutions.
@@
Conclusion (Null-space mechanism).
Under the linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
C.2 Reparameterization Symmetry (Existence Proof)
Setup: Neural networks exhibit inherent symmetries arising from overparameterization. We show these symmetries induce non-identifiability even in the exact nonlinear case.
Step 1: Representation reparameterization.
@@ C.2 Reparameterization Symmetry (Existence Proof)
Step 2: Linear reparameterizations.
@@
Note: The bias remains invariant under linear reparameterization through the origin. For more general affine transformations $T(h)=Ah+c$ with translation $c\neq 0$, the bias would also transform as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$. We restrict to origin-preserving transformations for simplicity, as these suffice to establish non-identifiability.
Then:
$h^{\prime}_{\ell+1}=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)=h_{\ell+1}.$ (25)
Step 3: Steering under reparameterization.
@@
$\tilde{h}^{\prime}_{\ell}=h^{\prime}_{\ell}+\alpha v^{\prime}=Ah_{\ell}+\alpha Av=A(h_{\ell}+\alpha v)=A\tilde{h}_{\ell}$. (28)
@@
Step 4: Infinitely many reparameterizations.
@@
Important note: While we cannot explicitly construct these reparameterizations for a frozen model without retraining (since this would require modifying $W_{\ell+1}$), their existence follows from the fundamental symmetry structure of neural networks. This establishes that identifiability cannot hold even in principle without additional constraints.
Practical implication: For frozen, deployed models, only the null-space mechanism is operationally relevant–the reparameterization symmetry is not a realizable operation. However, the reparameterization argument shows that even if we allowed weight modifications during training, identifiability would still fail without structural constraints. Thus, non-identifiability is both a practical limitation (null-space) and a fundamental theoretical barrier (gauge symmetry).
Conclusion (Reparameterization mechanism).
The inherent symmetries of neural network parameterizations induce infinitely many observationally equivalent steering vectors.
Appendix D Intuitive Explanations
D.1 Geometric Intuition for Null-Space Ambiguity
[…]
Figure 6: Geometric intuition for null-space ambiguity. The output observes only the $(v_{1},v_{2})$ components (blue shaded region). The $v_{3}$ component lies in $\ker(J)$ and is invisible to the output. Adding any $\alpha v_{3}$ to $v$ leaves the output unchanged: $J(v+\alpha v_{3})=Jv$ for all $\alpha\in\mathbb{R}$.
@@ D.1 Geometric Intuition for Null-Space Ambiguity
$J(v+\alpha v_{3})=Jv+\alpha J(0,0,1)^{\top}=Jv+\alpha\cdot 0=Jv.$ (30)
@@
D.2 Why More Data Doesn’t Help
@@
Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$. Each observation provides:
@@
D.3 Why Orthogonal Perturbations Preserve Semantics
@@
Observable effect is diffuse and unstructured.
@@ E.2 Fisher Information and Cramér-Rao Bound
$\mathcal{I}(v)=\mathbb{E}_{o}\left[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\right]=\frac{1}{\sigma^{2}}J^{\top}J.$ (42)
@@
Appendix F Disclosure of the Use of AI
We used ChatGPT to assist with refining sections of the manuscript for language and clarity and used Claude Code and GitHub Copilot for coding assistance during implementation. All research ideas, technical content, proofs, experiments and final writing decisions are entirely our own and we take full responsibility for the work presented.

### v3

@@ 6.4 Robustness of Non-Identifiability Under Distribution Shift
We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability. We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing) and safety-style (refusal patterns, ethical dilemmas). For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
Table 2 presents Cohen’s $d$ effect sizes across all trait-environment combinations for both models. Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$). Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across prompt contexts (in-dis: in-distribution, top: topic shift, gen: genre shift, saf: safety-style).
Model Trait In-dis Top Gen Saf
@@
Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models). Politeness exhibits higher variability ($d$ reaching 0.52–0.63 in genre-shift and safety contexts) but remains within the small-effect range. Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama). Figure 4 visualizes these patterns through environment-dependent effect sizes, showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
Figure 4: Environment-dependent effect sizes across prompt contexts.
Implications
The persistence of small effect sizes across diverse prompt distributions indicates that non-identifiability is not an artifact of in-distribution prompts but a structural geometric property. Cases where $d$ approaches 0.50.50.60.6 (e.g., politeness under genre or safety-style shifts) suggest that certain trait–context combinations impose partial constraints without eliminating null-space ambiguity since $\ker(J)$ is determined by model weights rather than the prompt distribution. Although effect sizes increase under distribution shift (reaching $d=0.63$ in some settings), distributional diversity alone does not eliminate observational equivalence.
7 Limitations
Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor). While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions. Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human evaluation, providing a conservative test of observational equivalence.
Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization or invariance objectives (IRM) can recover identifiable representations. Empirical validation of these approaches remains an important direction for future work. Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
8 Future Work
Our results establish the necessity of validation beyond input-output behavior. Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components and characterizing how null-space dimensionality scales with model size.
Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability. Human evaluation would provide stronger validation than our lexical heuristics and direct measurement of null-space dimensionality would quantify our theoretical predictions.
@@ 9 Conclusion
We have established that steering vectors are fundamentally non-identifiable under the local linear approximation, with infinitely many distinct directions producing equivalent behavioral effects. This non-identifiability persists across models, traits and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may fail to yield interpretable representations. Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone. More broadly, when internal interventions admit large equivalence classes under the model’s local geometry, consistent output shifts do not imply that a discovered direction corresponds to a canonical or causally privileged feature. Stronger interpretability claims therefore require structural constraints or additional assumptions that break these symmetries and distinguish meaningful factors from geometrically arbitrary ones.
Acknowledgments
We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments. We also acknowledge the HuggingFace platform and maintainers of the open-source models used in this work. Generative AI tools were used for language editing and code assistance; all technical content was reviewed and verified by the authors.
@@ Appendix A Logit-Level Equivalence Test
Values $R<1$ indicate that orthogonal perturbations induce smaller logit deviations than random directions. In all tested models and traits, this condition holds. Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions. The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$ and top-10 token overlap ranges from $86\%$ to $96\%$.
These results provide direct evidence that equivalence is not an artifact of semantic measurement. While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes. Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement and top-$k$ overlap for all models and traits.
Table 3: Logit-level equivalence statistics. $R$ denotes the ratio $v_{\perp}$ / random. Agreement% refers to token agreement, and Overlap% indicates top-10 overlap.
Model Trait $R$ Agreement% Overlap%
@@
Formality 0.76 94% 96%
Politeness 0.75 93% 96%
@@
Formality 0.49 93% 95%
Politeness 0.65 91% 94%
Appendix B Detailed Proof of Proposition 1
B.1 Null-Space Ambiguity (Constructive Proof)
Setup Consider the local linear approximation of the steering effect:
@@ B.1 Null-Space Ambiguity (Constructive Proof)
Step 1: Null space characterization
@@
Step 2: Rank Bound
@@
In modern language models, the hidden dimension is typically $d\approx 4000$ while the vocabulary size is $V\approx 50000$. Therefore, the Jacobian satisfies $\max\mathrm{rank}(J_{\ell})=d$.
Step 3: Effective rank is much lower
@@
Intuition: Empirical evidence suggests that in overparameterized LLMs the effective rank of $J_{\ell}$ may be substantially lower than $d$, consistent with low dimensional structure observed in related weight space studies (Li et al., 2018; Maennel et al., 2018) and concentration of output sensitivity in a limited number of directions (Ghorbani et al., 2019). Consequently, $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is often positive in practice. However, this empirical observation is not required for the proof; the argument establishes non identifiability whenever $\dim(\mathcal{N})\geq 1$.
Step 4: Constructing equivalent vectors
@@
$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (18)
$=o(x,0,0)+\alpha J_{\ell}(x)v$
$\approx o(x,v,\alpha).$
Step 5: Infinitely many distinct solutions
@@
Conclusion (Null-space mechanism)
Under the local linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
B.2 Weight Space Gauge Symmetry (Supplementary Observation)
While Proposition 1 fixes model weights and asks whether v is recoverable from behavior, this section offers a complementary perspective: even relaxing the weight-fixing assumption, non-identifiability persists via linear reparameterization symmetries.
Setup: We record a supplementary observation about gauge symmetry in neural network parameterization. This is distinct from the null-space non-identifiability of Proposition 1: rather than fixing model weights and asking whether $v$ is recoverable from behavioral observations, we ask whether different $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs can produce identical input–output behavior. The answer is yes, via the linear reparameterization below. This symmetry is not realizable as a steering operation on a frozen model as it requires modifying $W_{\ell+1}$.
Step 1: Representation reparameterization
@@ B.2 Weight Space Gauge Symmetry (Supplementary Observation)
Step 2: Linear reparameterizations
@@
Note: The bias remains invariant under linear reparameterization through the origin. For more general affine transformations $T(h)=Ah+c$ with translation $c\neq 0$, the bias would also transform as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$. We restrict to origin-preserving transformations for simplicity, as these suffice to establish non-identifiability. Then:
$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (25)
$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
$=h_{\ell+1}.$
Step 3: Steering under reparameterization
@@
$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (28)
$=Ah_{\ell}+\alpha Av$
$=A(h_{\ell}+\alpha v)$
$=A\tilde{h}_{\ell}$.
@@
Step 4: Infinitely many reparameterizations
@@
Conclusion (Weight-space gauge symmetry)
The symmetries of neural network parameterization imply the existence of infinitely many $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs that yield identical input–output behavior.
Appendix C Intuitive Explanations
C.1 Geometric Intuition for Null-Space Ambiguity
@@
[…]
Figure 5: Geometric intuition for null-space ambiguity. The output depends only on the $(v_{1},v_{2})$ projection (row space of $J$), while $v_{3}\in\ker(J)$ is invisible; hence $J(v+\alpha v_{3})=Jv$.
@@
$J(v+\alpha v_{3})$ $=Jv+\alpha Jv_{3}$ (30)
$=Jv+\alpha\cdot 0\quad(v_{3}\in\ker(J))$
$=Jv.$
@@
C.2 Why More Data Doesn’t Help
@@
Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$. More fundamentally, any $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v_{0}=0$ by definition since this is a property of the model weights and not the prompt distribution. Stacking observations across prompts therefore cannot remove directions from $\ker(J_{\ell})$. Each observation provides:
@@
C.3 Why Orthogonal Perturbations Preserve Semantics
@@
Observable effect is diffuse and unstructured
@@
Appendix D Prompt Construction Details
Table 4: Example contrastive prompt pairs for each semantic trait.
Trait Positive ($x^{+}$) Negative ($x^{-}$)
Figure 7: Steering extraction and evaluation pipeline.
@@ E.2 Fisher Information and Cramér-Rao Bound
$\mathcal{I}(v)$ $=\mathbb{E}_{o}\Bigg[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)$ (42)
$\qquad\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\Bigg]$
$=\frac{1}{\sigma^{2}}J^{\top}J.$